# Spark Transformer: Reactivating Sparsity in FFN and Attention

### v1

@@ 2.2 Spark Attention
where we omit the dependency on $i$ for simplicity. The computation cost associated with Equation 6 is $4d_{\text{attn}}\cdot n_{\text{ctx}}$ for each head. Finally, output from all heads are concatenated followed by a linear map to project to $d_{\text{model}}$.
Note that Equation 6 has the same form as FFN in Equation 1 except for the choice of nonlinearity. Hence, following a similar strategy in obtaining Spark FFN, here we present Spark Attention as
[…]
In above, ${\bm{K}}_{1}\in\mathbb{R}^{r\times n_{\text{ctx}}}$ and ${\bm{K}}_{2}\in\mathbb{R}^{(d_{\text{attn}}-r)\times n_{\text{ctx}}}$ contain the first $r$ rows and the rest of the rows from ${\bm{K}}$, respectively, and $\sigma_{1}$ and $\sigma_{2}$ are nonlinear functions. Empirically, we find that taking $\sigma_{1}=\operatorname{softmax}$ and $\sigma_{2}=\operatorname{softplus}$ gives good results. Finally, $\operatorname{Top}_{k}^{(-\infty)}$ refers to an operator that keeps the largest $k$ values of the input while setting the rest to $\infty$- ∞. This operator be implemented using statistical top-k, which we explain in Section 3.
FLOPs per Token. With a direct implementation the number of FLOPs in Equation 7 is given by $4d_{\text{attn}}\cdot n_{\text{ctx}}$, which is the same as the FLOPs for Equation 6. However, by noting that the output of the softmax is expected to be sparse with approximately $k$ nonzero entries, the computation costs associated with ${\bm{K}}_{1}^{\top}{\bm{q}}[:\!r]$ and in the multiplication with ${\bm{V}}$ can be drastically reduced. In particular, if we take $r=\frac{d_{\text{attn}}}{2}$ then the FLOPs per token becomes
@@
which is nearly a 4×4$\times$4 × reduction when $k_{\text{attn}}\ll n_{\text{ctx}}$.
@@ 3 Statistical Top-k
This section introduces $\operatorname{{Statistical-Top}}_{k}$, an efficient algorithm for implementing the $\operatorname{Top}_{k}$ operators in Spark FFN (i.e. Equation 2) and Spark Attention (i.e. Equation 7).
@@
where ${\bm{1}}$ and ${\bm{0}}$ are $d$-dimensional vectors with all entries equal to $1$ and 00, respectively. The soft-thresholding operator shifts each entry of ${\bm{x}}$ to the left by $\theta$ and then thresholds the result at zero.
@@
In above, we define $\operatorname{mean}({\bm{x}})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{d}\sum_{i=1}^{d}x_{i}$ and $\operatorname{std}({\bm{x}})\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{\frac{1}{d-1}\sum_{i=1}^{d}(x_{i}-\operatorname{mean}({\bm{x}}))^{2}}$ , which compute the sample mean and standard deviation of the entries of the input ${\bm{x}}$, respectively. $Q(\cdot)$ is the quantile function (i.e., inverse of the cumulative distribution function) of the standard Gaussian distribution. In Spark Transformer, Equation 10 is used as the $\operatorname{Top}_{k}$ operator in Equation 2. For the operator $\operatorname{Top}_{k}^{(-\infty)}$ in Equation 7, a slight variant of Equation 10 is used where the entries below the threshold $\theta({\bm{x}},k)$ are set to $\infty$- ∞ instead of 00.
@@ 3.1 Threshold Estimation
The threshold $\theta({\bm{x}},k)$ in Equation 10 is designed such that, if the entries of ${\bm{x}}$ are drawn from a Gaussian distribution, approximately $k$ out of the $d$ entries will exceed this threshold. To understand this, let $\mu$ and $\sigma$ denote the mean and standard deviation of the underlying Gaussian distribution. Its quantile function is given by $\mu+\sigma\cdot Q(p)$ for $p\in(0,1)$. Consequently, due to the properties of quantile functions, we expect roughly $p\cdot d$ entries of ${\bm{x}}$ to exceed $\mu+\sigma\cdot Q(1-p)$. In practice, since $\mu$ and $\sigma$ are unknown, they are replaced with the sample mean $\operatorname{mean}({\bm{x}})$ and the sample standard deviation $\operatorname{std}({\bm{x}})$, respectively.
@@
Let ${\bm{x}}\in\mathbb{R}^{d}$ be a vector with entries drawn i.i.d. from $\mathcal{N}(\mu,\sigma^{2})$. For any $11\leq k\leq d-$ 1, let $\theta({\bm{x}},k)$ be a scalar defined in Equation 10. Take any $\delta\in(0,1)$ and assume $d\geq\max\{2,\log\frac{6}{\delta}\}$. With a probability of at least $1-\delta$, the number of entries of ${\bm{x}}$ that are greater than $\theta({\bm{x}},k)$, i.e., $\operatorname{\mathbf{card}}\left(\{i\in[d]~{}|~{}x_{i}>\theta({\bm{x}},k)\}\right)$, satisfies
$\frac{\left|\operatorname{\mathbf{card}}\left(\{i\in[d]~{}|~{}x_{i}>\theta({\bm{x}},k)\}\right)-k\right|}{d}\leq 4\sqrt{\frac{\log\frac{6}{\delta}}{d}}\left(1+\sqrt{-2\log\min\left\{\frac{k}{d},1-\frac{k}{d}\right\}}\right)\,.$ (12)
Theorem 3.1 provides a relative error bound between $k$ and the true number of entries of ${\bm{x}}$ that exceed $k$. This bound is maximized when $k=1$ or $k=d-1$. Consequently, the worst-case bound is $O\left(\sqrt{\frac{\log d\cdot\log\frac{1}{\delta}}{d}}\right)$ which vanishes as $d$ increases. Notably, the error bound becomes $O\left(\sqrt{\frac{\log\frac{1}{\delta}}{d}}\right)$ when $k=\Theta(d)$, demonstrating even faster convergence.
@@
While the Gaussian quantile function $Q(\cdot)$ lacks a closed-form solution, high-precision piecewise approximation algorithms with constant complexity are available in standard software packages like SciPy [81], readily applicable to our needs.
@@ 3.2 Sparsification
Given the threshold $\theta({\bm{x}},k)$, a straightforward approach to obtain a sparse vector is to set all entries of ${\bm{x}}$ below the threshold to zero, preserving the remaining values. This operator, sometimes referred to as hard thresholding [10], suffers from discontinuity, potentially hindering its suitability for gradient-descent-based training.
To address this, $\operatorname{{Statistical-Top}}_{k}$ employs the soft-thresholding operator defined in Equation 9 [8]. This operator first shrinks all entries of ${\bm{x}}$ by the threshold $\theta({\bm{x}},k)$ and then sets all entries below 00 to 00. Soft thresholding offers the advantages of being continuous and differentiable almost everywhere (except when entries of ${\bm{x}}$ coincide with $\theta({\bm{x}},k)$).
For complete differentiability, one can utilize a smoothing function like the Huber loss [35], defined element-wise on an input ${\bm{x}}$ as:
$\operatorname{Huber}(x;\delta)\stackrel{{\scriptstyle\text{def}}}{{=}}\begin{cases}\frac{1}{2}x^{2}&\text{for}~{}|x|<\delta,\\ \delta\cdot(|x|-\frac{1}{2}\delta)&\text{otherwise}.\end{cases}$ (13)
@@
Note that Equation 14 converges to $\operatorname{{Statistical-Top}}_{k}({\bm{x}})$ as $\delta\to 0$, which can be seen as $\operatorname{Huber}(x;\delta)/\delta\to|x|$ and $\operatorname{{Statistical-Top}}_{k}({\bm{x}})$ is always non-negative. In practice, however, we find that using a non-zero $\delta$ does not improve model quality, and therefore we set $\delta=0$ for simplicity.
Finally, soft thresholding admits a variational form (see, e.g., [62]):
@@
This formulation seeks a vector $\mathbf{z}$ that minimizes both its squared $\ell_{2}$ distance to the input ${\bm{x}}$ and its $\ell_{1}$ norm, with the threshold $\theta$ balancing these terms. Given the sparsity-promoting nature of the $\ell_{1}$ norm, soft thresholding effectively finds a sparse approximation of the input ${\bm{x}}$. This variational form also connects $\operatorname{{Statistical-Top}}_{k}$ with other top-$k$ algorithms in the literature; see Section 3.3.
@@ 3.3 Comparison of Statistical Top-$k$ with Related Top-$k$ Operators
The variational form in Equation 15 reveals connections of $\operatorname{{Statistical-Top}}_{k}$ with other top-$k$ algorithms in the literature. Specifically, [47] defines a soft top-$k$ as
$\operatorname*{arg\,min}_{{\bm{z}}}~{}~{}-\theta\cdot H({\bm{z}})-\langle{\bm{z}},{\bm{x}}\rangle,~{}~{}\text{s.t.}~{}~{}{\bm{z}}^{\top}{\bm{1}}=k,~{}~{}{\bm{0}}\leq{\bm{z}}\leq{\bm{1}},$ (16)
where $H({\bm{z}})$ is the entropy function. Another work [54] defines the SparseK operator
$\operatorname*{arg\,min}_{{\bm{z}}}~{}~{}-H^{G}({\bm{z}})-\langle{\bm{z}},{\bm{x}}\rangle,~{}~{}\text{s.t.}~{}~{}{\bm{z}}^{\top}{\bm{1}}=k,~{}~{}{\bm{0}}\leq{\bm{z}}\leq{\bm{1}},$ (17)
@@
$\operatorname{{Statistical-Top}}_{k}$ in the form of Equation 15, as well as Equation 16 and Equation 17, can all be interpreted as finding an output that is close to the input subject to a sparsifying regularization. Their major difference lies in the choice of the sparse regularization. That is, soft top-$k$ and SparseK uses entropy and Gini entropy, respectively, whereas $\operatorname{{Statistical-Top}}_{k}$ uses $\ell_{1}$. The choice of $\ell_{1}$ makes $\operatorname{{Statistical-Top}}_{k}$ superior in that it has a closed form solution provided by soft-thresholding, which only requires $d$ FLOPs. In contrast, soft top-$k$ and SparseK both do not have closed form solutions and require an iterative algorithm with a FLOP count dependent on the number of iterations. In addition, there is no guarantee that soft top-$k$ and SparseK can obtain (approximately) $k$ nonzero entries as output.
Finally, we mention that ideas similar to statistical top-$k$ have been used [74, 57] for the problem of distributed training [50]. However, we are the first to introduce, adapt, and verify its effectiveness for activation sparsity. Additional discussions are provided in Section D.2.
@@ 4 Experiments
In this section, we present an experimental evaluation of Spark Transformer using the Gemma-2 2B model. Gemma-2 2B is a decoder-only Transformer with 2 billion parameters, pretrained on 2 trillion tokens of primarily English text data (see [27] for details). To evaluate Spark Transformer, we train a model by substituting the standard FFN and Attention in Gemma-2 2B with their Spark Transformer counterparts (Spark FFN and Spark Attention, respectively). This Spark Transformer model is trained using the same procedure and data as the Gemma-2 2B model.
Implementation details. Gemma-2 uses a model dimension of $d_{\text{model}}=2304$. For FFN, Gemma-2 uses the Gated FFN in Equation 5 with $d_{\text{ff}}^{\prime}=9216$. We replace it with Spark FFN in Equation 2 with $d_{\text{ff}}=13824$ so that the parameter count keeps the same. In addition, we take $k$ to be 1106, which gives a sparsity level of $8\%$, and $r=1024\approx d_{\text{model}}/2$ (due to sharding constraints, $r$ can only be a multiple of $256$). For Attention, Gemma-2 alternates between a global attention that have a span of 8192 tokens, and a local attention with a 4096 window size, both with $d_{\text{attn}}=256$. We replace both with Spark Attention in Equation 6 where for the latter we use the same 4096 window size. For hyper-parameters, we use $k=256$, i.e. each token attends to at most $256$ tokens, and $r=128=d_{\text{attn}}/2$. Gemma-2 uses Rotary Position Embedding [79] which is applied to ${\bm{q}}$ and the columns of ${\bm{K}}$ in Equation 6. For Spark Attention in Equation 7, we apply this position encoding to ${\bm{q}}[:\!r]$, ${\bm{q}}[r\!:]$, the columns of ${\bm{K}}_{1}$, and the columns of ${\bm{K}}_{2}$.
@@
Figure 3: Sparsity in the intermediate activation of Spark FFN and Spark Attention across $26$ layers at selected training steps. For FFN we report the percentage of nonzero entries out of $d_{\text{ff}}=13824$ entries. For Attention, we report the number of nonzero entries (i.e., attended tokens). Our hyper-parameter choice is to have 8% nonzeros in Spark FFN and at most 256 nonzeros in Spark Attention.

### v2

@@ 2.2 Spark Attention
where we omit the dependency on $i$ for simplicity. The computation cost associated with Equation˜6 is $4d_{\text{attn}}\cdot n_{\text{ctx}}$ for each head. Finally, output from all heads are concatenated followed by a linear map to project to $d_{\text{model}}$.
Note that Equation˜6 has the same form as FFN in Equation˜1 except for the choice of nonlinearity. Hence, following a similar strategy in obtaining Spark FFN, here we present Spark Attention as
$\text{Spark-Attention}({\bm{q}};{\bm{K}},{\bm{V}},k,r)\stackrel{{\scriptstyle\text{def}}}{{=}}{\bm{V}}\cdot\bigg(\sigma_{1}\!\left(\operatorname{Top}_{k}^{(-\infty)}({\bm{K}}_{1}^{\top}{\bm{q}}[:\!r])\right)\odot\sigma_{2}\!\left({\bm{K}}_{2}^{\top}{\bm{q}}[r\!:]\right)\bigg).$ (7)
In above, ${\bm{K}}_{1}\in\mathbb{R}^{r\times n_{\text{ctx}}}$ and ${\bm{K}}_{2}\in\mathbb{R}^{(d_{\text{attn}}-r)\times n_{\text{ctx}}}$ contain the first $r$ rows and the rest of the rows from $\bm{K}$}, respectively, and $\sigma_{1}$ and $\sigma_{2}$ are nonlinear functions. Empirically, we find that taking $\sigma_{1}=\operatorname{softmax}$ and $\sigma_{2}=\operatorname{softplus}$ gives good results. Finally, $\operatorname{Top}_{k}^{(-\infty)}$ refers to an operator that keeps the largest $k$ values of the input while setting the rest to -$\infty$. This operator be implemented using statistical top-k, which we explain in Section˜3.
FLOPs per Token. With a direct implementation the number of FLOPs in Equation˜7 is given by $4d_{\text{attn}}\cdot n_{\text{ctx}}$, which is the same as the FLOPs for Equation˜6. However, by noting that the output of the softmax is expected to be sparse with approximately $k$ nonzero entries, the computation costs associated with ${\bm{K}}_{1}^{\top}{\bm{q}}[:\!r]$ and in the multiplication with $\bm{V}$} can be drastically reduced. In particular, if we take $r=\frac{d_{\text{attn}}}{2}$ then the FLOPs per token becomes
@@
which is nearly a $4\times$ reduction when $k_{\text{attn}}\ll n_{\text{ctx}}$.
@@ 3 Statistical Top-k
This section introduces $\operatorname{{Statistical-Top}}_{k}$, an efficient algorithm for implementing the $\operatorname{Top}_{k}$ operators in Spark FFN (i.e. Equation˜2) and Spark Attention (i.e. Equation˜7).
@@
where $\bm{1}$} and $\bm{0}$} are $d$-dimensional vectors with all entries equal to $1$ and 0, respectively. The soft-thresholding operator shifts each entry of $\bm{x}$} to the left by $\theta$ and then thresholds the result at zero.
@@
In above, we define $\operatorname{mean}({\bm{x}})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{d}\sum_{i=1}^{d}x_{i}$ and $\operatorname{std}({\bm{x}})\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{\frac{1}{d-1}\sum_{i=1}^{d}(x_{i}-\operatorname{mean}({\bm{x}}))^{2}}$ , which compute the sample mean and standard deviation of the entries of the input $\bm{x}$}, respectively. $Q(\cdot)$ is the quantile function (i.e., inverse of the cumulative distribution function) of the standard Gaussian distribution. In Spark Transformer, Equation˜10 is used as the $\operatorname{Top}_{k}$ operator in Equation˜2. For the operator $\operatorname{Top}_{k}^{(-\infty)}$ in Equation˜7, a slight variant of Equation˜10 is used where the entries below the threshold $\theta({\bm{x}},k)$ are set to -$\infty$ instead of 0.
@@ 3.1 Threshold Estimation
The threshold $\theta({\bm{x}},k)$ in Equation˜10 is designed such that, if the entries of $\bm{x}$} are drawn from a Gaussian distribution, approximately $k$ out of the $d$ entries will exceed this threshold. To understand this, let $\mu$ and $\sigma$ denote the mean and standard deviation of the underlying Gaussian distribution. Its quantile function is given by $\mu+\sigma\cdot Q(p)$ for $p\in(0,1)$. Consequently, due to the properties of quantile functions, we expect roughly $p\cdot d$ entries of $\bm{x}$} to exceed $\mu+\sigma\cdot Q(1-p)$. In practice, since $\mu$ and $\sigma$ are unknown, they are replaced with the sample mean $\operatorname{mean}({\bm{x}})$ and the sample standard deviation $\operatorname{std}({\bm{x}})$, respectively.
@@
Let ${\bm{x}}\in\mathbb{R}^{d}$ be a vector with entries drawn i.i.d. from $\mathcal{N}(\mu,\sigma^{2})$. For any $1\leq k\leq d-1$, let $\theta({\bm{x}},k)$ be a scalar defined in Equation˜10. Take any $\delta\in(0,1)$ and assume $d\geq\max\{2,\log\frac{6}{\delta}\}$. With a probability of at least $1-\delta$, the number of entries of $\bm{x}$} that are greater than $\theta({\bm{x}},k)$, i.e., $\operatorname{\mathbf{card}}\left(\{i\in[d]\penalty 10000\ |\penalty 10000\ x_{i}>\theta({\bm{x}},k)\}\right)$, satisfies
$\frac{\left|\operatorname{\mathbf{card}}\left(\{i\in[d]\penalty 10000\ |\penalty 10000\ x_{i}>\theta({\bm{x}},k)\}\right)-k\right|}{d}\leq 4\sqrt{\frac{\log\frac{6}{\delta}}{d}}\left(1+\sqrt{-2\log\min\left\{\frac{k}{d},1-\frac{k}{d}\right\}}\right)\,.$ (12)
Theorem 3.1 provides a relative error bound between $k$ and the true number of entries of $\bm{x}$} that exceed $k$. This bound is maximized when $k=1$ or k=d1k=d-1. Consequently, the worst-case bound is $O\left(\sqrt{\frac{\log d\cdot\log\frac{1}{\delta}}{d}}\right)$ which vanishes as $d$ increases. Notably, the error bound becomes $O\left(\sqrt{\frac{\log\frac{1}{\delta}}{d}}\right)$ when $k=\Theta(d)$, demonstrating even faster convergence.
@@
While the Gaussian quantile function $Q(\cdot)$ lacks a closed-form solution, high-precision piecewise approximation algorithms with constant complexity are available in standard software packages like SciPy (Virtanen et al., 2020), readily applicable to our needs.
@@ 3.2 Sparsification
Given the threshold $\theta({\bm{x}},k)$, a straightforward approach to obtain a sparse vector is to set all entries of $\bm{x}$} below the threshold to zero, preserving the remaining values. This operator, sometimes referred to as hard thresholding (Blumensath and Davies, 2008), suffers from discontinuity, potentially hindering its suitability for gradient-descent-based training.
To address this, $\operatorname{{Statistical-Top}}_{k}$ employs the soft-thresholding operator defined in Equation˜9 (Beck and Teboulle, 2009). This operator first shrinks all entries of $\bm{x}$} by the threshold $\theta({\bm{x}},k)$ and then sets all entries below 0 to 0. Soft thresholding offers the advantages of being continuous and differentiable almost everywhere (except when entries of $\bm{x}$} coincide with $\theta({\bm{x}},k)$).
For complete differentiability, one can utilize a smoothing function like the Huber loss (Huber, 1992), defined element-wise on an input $\bm{x}$} as:
$\operatorname{Huber}(x;\delta)\stackrel{{\scriptstyle\text{def}}}{{=}}\begin{cases}\frac{1}{2}x^{2}&\text{for}\penalty 10000\ |x|<\delta,\\ \delta\cdot(|x|-\frac{1}{2}\delta)&\text{otherwise}.\end{cases}$ (13)
@@
Note that Equation˜14 converges to $\operatorname{{Statistical-Top}}_{k}({\bm{x}})$ as $\delta\to 0$, which can be seen as $\operatorname{Huber}(x;\delta)/\delta\to|x|$ and $\operatorname{{Statistical-Top}}_{k}({\bm{x}})$ is always non-negative. In practice, however, we find that using a non-zero $\delta$ does not improve model quality, and therefore we set $\delta=0$ for simplicity.
Finally, soft thresholding admits a variational form (see, e.g., Parikh et al. (2014)):
@@
This formulation seeks a vector $\mathbf{z}$ that minimizes both its squared $\ell_{2}$ distance to the input $\bm{x}$} and its $\ell_{1}$ norm, with the threshold $\theta$ balancing these terms. Given the sparsity-promoting nature of the $\ell_{1}$ norm, soft thresholding effectively finds a sparse approximation of the input $\bm{x}$}. This variational form also connects $\operatorname{{Statistical-Top}}_{k}$ with other top-$k$ algorithms in the literature; see Section˜3.3.
@@ 3.3 Comparison of Statistical Top-$k$ with Related Top-$k$ Operators
The variational form in Equation˜15 reveals connections of $\operatorname{{Statistical-Top}}_{k}$ with other top-$k$ algorithms in the literature. Specifically, Lei et al. (2023) defines a soft top-$k$ as
$\operatorname*{arg\,min}_{{\bm{z}}}\penalty 10000\ \penalty 10000\ -\theta\cdot H({\bm{z}})-\langle{\bm{z}},{\bm{x}}\rangle,\penalty 10000\ \penalty 10000\ \text{s.t.}\penalty 10000\ \penalty 10000\ {\bm{z}}^{\top}{\bm{1}}=k,\penalty 10000\ \penalty 10000\ {\bm{0}}\leq{\bm{z}}\leq{\bm{1}},$ (16)
where $H({\bm{z}})$ is the entropy function. Another work (Lou et al., 2024) defines the SparseK operator
$\operatorname*{arg\,min}_{{\bm{z}}}\penalty 10000\ \penalty 10000\ -H^{G}({\bm{z}})-\langle{\bm{z}},{\bm{x}}\rangle,\penalty 10000\ \penalty 10000\ \text{s.t.}\penalty 10000\ \penalty 10000\ {\bm{z}}^{\top}{\bm{1}}=k,\penalty 10000\ \penalty 10000\ {\bm{0}}\leq{\bm{z}}\leq{\bm{1}},$ (17)
@@
$\operatorname{{Statistical-Top}}_{k}$ in the form of Equation˜15, as well as Equation˜16 and Equation˜17, can all be interpreted as finding an output that is close to the input subject to a sparsifying regularization. Their major difference lies in the choice of the sparse regularization. That is, soft top-$k$ and SparseK uses entropy and Gini entropy, respectively, whereas $\operatorname{{Statistical-Top}}_{k}$ uses $\ell_{1}$. The choice of $\ell_{1}$ makes $\operatorname{{Statistical-Top}}_{k}$ superior in that it has a closed form solution provided by soft-thresholding, which only requires $d$ FLOPs. In contrast, soft top-$k$ and SparseK both do not have closed form solutions and require an iterative algorithm with a FLOP count dependent on the number of iterations. In addition, there is no guarantee that soft top-$k$ and SparseK can obtain (approximately) $k$ nonzero entries as output.
Finally, we mention that ideas similar to statistical top-$k$ have been used (Shi et al., 2019; M Abdelmoniem et al., 2021) for the problem of distributed training (Lin et al., 2018). However, we are the first to introduce, adapt, and verify its effectiveness for activation sparsity. Additional discussions are provided in Section˜D.2.
@@ 4 Experiments
In this section, we present an experimental evaluation of Spark Transformer using the Gemma-2 2B model. Gemma-2 2B is a decoder-only Transformer with 2 billion parameters, pretrained on 2 trillion tokens of primarily English text data (see Gemma Team (2024) for details). To evaluate Spark Transformer, we train a model by substituting the standard FFN and Attention in Gemma-2 2B with their Spark Transformer counterparts (Spark FFN and Spark Attention, respectively). This Spark Transformer model is trained using the same procedure and data as the Gemma-2 2B model.
Implementation details. Gemma-2 uses a model dimension of $d_{\text{model}}=2304$. For FFN, Gemma-2 uses the Gated FFN in Equation˜5 with $d_{\text{ff}}^{\prime}=9216$. We replace it with Spark FFN in Equation˜2 with $d_{\text{ff}}=13824$ so that the parameter count keeps the same. In addition, we take $k$ to be 1106, which gives a sparsity level of $8\%$, and $r=1024\approx d_{\text{model}}/2$ (due to sharding constraints, $r$ can only be a multiple of $256$). For Attention, Gemma-2 alternates between a global attention that have a span of 8192 tokens, and a local attention with a 4096 window size, both with $d_{\text{attn}}=256$. We replace both with Spark Attention in Equation˜6 where for the latter we use the same 4096 window size. For hyper-parameters, we use $k=256$, i.e. each token attends to at most $256$ tokens, and $r=128=d_{\text{attn}}/2$. Gemma-2 uses Rotary Position Embedding (Su et al., 2024) which is applied to $\bm{q}$} and the columns of $\bm{K}$} in Equation˜6. For Spark Attention in Equation˜7, we apply this position encoding to ${\bm{q}}[:\!r]$, ${\bm{q}}[r\!:]$, the columns of ${\bm{K}}_{1}$, and the columns of ${\bm{K}}_{2}$.
@@
Figure 3: Sparsity in the intermediate activation of Spark FFN and Spark Attention across 2626 layers at selected training steps. For FFN we report the percentage of nonzero entries out of $d_{\text{ff}}=13824$ entries. For Attention, we report the number of nonzero entries (i.e., attended tokens). Our hyper-parameter choice is to have 8% nonzeros in Spark FFN and at most 256 nonzeros in Spark Attention.